# A Comprehensive Survey on Network Traffic Synthesis: From Statistical Models to Deep Learning

### v1

@@ V-C4 Light-weight DMs
Another angle of improvement for DMs were addressed in making the model lightweight such that it can be used for high-quality data generation at edge devices. [83] proposes a DM specifically designed for resource-constrained edge devices. By integrating depthwise separable convolutions into the diffusion process, this model achieves significant reductions in computational overhead while maintaining the diversity and quality of generated malicious traffic. The evaluations demonstrate the ability of the model to produce high-quality synthetic data, improving the training of intrusion detection classifiers even on hardware-limited edge nodes. This lightweight design underscores the potential of diffusion models to address computational challenges, enabling real-time cybersecurity applications at the edge.
V-C5 Limitations
@@
V-C6 Summary
The emergence of DMs has revolutionized the field of network traffic analysis, providing a powerful alternative to traditional generative models like GANs. Table VI provides a comprehensive summary of the work done with DMs. Innovative adaptations of DMs, such as DDPM, stable diffusion models, and controlled diffusion models, have demonstrated remarkable capabilities to generate realistic high-quality network traffic data. By effectively modeling the underlying distributions of packet-level features and user interactions with network services, DMs have consistently outperformed previous SOTA GANs in terms of accuracy and realism. The ability of DMs to capture complex temporal and feature correlations while mitigating issues such as mode collapse has solidified their position as the leading generative framework for network traffic synthesis. Further, lightweight diffusion models make the way for high-quality data synthesis in a resource-constrained environment. As research in this area continues to advance, DMs are poised to play an increasingly crucial role in enhancing network security, improving data analysis, and addressing the challenges of data scarcity in network traffic environments.
V-D Transformers
@@
Bikmukhamedov et al.[18] Generic network traffic generation, Attack detection GPT-2 Transport, Network Packet level and Flow level No
Meng et al.[99] Generic network traffic generation GPT-2 Transport, Network Packet level and Flow level No
Qu et al.[124] Network optimization and Long traffic flow classification GPT + Linear attention mechanism Network Flow level No
Kholgh et al.[71] Generic network traffic generation, IDS GPT-3 Transport, Network Packet level, Flow level No
Kong et al.[77] Cellular network traffic synthesis CPT-GPT Application Event-level sequences No
Transformers were first introduced in 2017 in [152] to address sequence transduction tasks, such as neural machine translation. These tasks involve converting an input sequence into an output sequence. The transformer model is a type of neural network that learns the context of sequential data and generates new data based on its understanding of the relationships between elements within the sequence. Transformers excel at understanding context by using a mechanism called attention, which enables the model to weigh the importance of different elements in a sequence.
@@
V-D1 Generative Pretrained Transformers (GPT)
@@
With the introduction of GPT-3 models, authors of  [71] leveraged GPT-3 Davinci model and GPT-3 Babbage model to synthesize network traffic for various tasks. The pipeline of this process would be triggered by a user requesting network traffic either using a specific scenario (e.g., normal traffic data, a Ping-of-Death attack, etc.) or by specifying a set of network protocols (e.g., ICMP, DNS, etc.) Then, a Flow Generator, which is a Python script, produces network flow in text format, i.e. it generates the sequence and textual summary of packets that make up the network traffic, although not creating the actual packets themselves. The flow summary is then passed to the Packet Generator, which is a transformer-based model capable of creating packets using the network flow. The packet generator then creates the specified packets and writes them in .pcap format to a file. GPT-3 Davinci was used to generate training samples to train the model as it was a more accurate GPT-3 model. Then, to reduce the cost and to get faster responses while inferencing, GPT-3 Babbage model was used in the packet generator.
Similarly,  [77] presents CPT-GPT, a high-fidelity traffic generator tailored for cellular network control-plane traffic. CPT-GPT effectively addresses challenges such as stateful semantic correctness, multimodal feature relationships, and long-term data drifts by leveraging a Transformer-based architecture. Through its innovative tokenization scheme and probabilistic output mechanism, CPT-GPT generates realistic traffic while avoiding issues like state violations and mode collapse. Unlike traditional Semi-Markov models (SMMs) that require extensive manual updates to align with evolving 3GPP standards, CPT-GPT autonomously achieves comparable or superior fidelity.
V-D2 Limitations
@@ V-D3 Summary
The integration of transformer models, particularly GPT, into the realm of network traffic generation represents a significant advancement in the field of artificial intelligence and data synthesis. Table VII provides a comprehensive summary of the work done with Transformers. The evolution from initial applications in natural language processing to specialized frameworks like NetGPT and TrafficGPT showcases the versatility of transformers in handling complex data structures, such as network traffic. Innovations in tokenization and attention mechanisms have addressed challenges like flow length limitations and class imbalance, leading to improved performance in synthetic traffic generation. As these models continue to evolve, they hold the promise of transforming network analysis, enhancing security measures, and providing valuable insights into network behavior through more accurate and comprehensive traffic generation methodologies.
Table VIII: Comparative analysis of synthetic network traffic generation approaches (considers only the vanilla models)
Aspect Statistical Modeling Deep Learning–Based Methods
VAEs GANs Diffusion Models Transformers
Data Requirements Moderate: tuned from real measurements, fewer data required High: requires sufficient data to learn a meaningful latent space Very High: demands large volumes to avoid mode collapse and capture diversity Very High: extensive data needed to capture fine–grained, complex details Very High: typically trained on massive corpora to learn long–term dependencies
Model Complexity Low: based on distributions and parameter estimation Moderate: encoder-decoder structure with multiple layers Moderate: requires more training, sensitive to hyperparameters and mode collapse Very High: iterative denoising with multi–step sampling increases complexity High: utilizes self–attention layers with large scale pretraining
Training Stability High: analytical models yield deterministic solutions Low: stability can be affected by the trade–off between reconstruction and regularization Low: adversarial training is inherently unstable and sensitive to hyperparameters causing mode collapse Moderate: generally stable training Moderate: with proper optimization, training can be stable
Capture Complex Patterns Low: effective for basic statistical properties but struggles with non–linear dynamics Moderate: can capture temporal features, especially when enhanced with sequential layers High: well–suited for learning complex, non–linear patterns (long–term dependencies can be challenging) Very High: excels at capturing both fine–grained and long-range temporal dependencies Very High: particularly adept at modeling long–term dependencies and sequential structures
Interpretability High: parameters and distributions have clear statistical meanings Low: latent space analysis is needed for interpretability Low: adversarial frameworks provide limited insight into internal representations Low: the iterative denoising process is less intuitive Moderate: attention mechanisms offer partial insight despite overall complexity
Computational Efficiency High: typically lightweight and fast due to closed–form calculations Moderate: inference is efficient, but training can be resource–intensive Moderate: training two networks simultaneously increases computational cost Low: iterative sampling during inference is computationally expensive Low: training large transformer models (GPT) requires significant resources
Data Fidelity Moderate: capable of reproducing basic traffic patterns, but may not capture high variability or complex interactions Moderate: produces plausible outputs but may appear over–smoothed or lack fine details High: when well–trained, generates highly realistic traffic that closely mimics real–world nuances Very High: achieves SOTA fidelity in replicating complex patterns and fine details Very High: produces coherent, context–aware sequences reflecting long–term dependencies
Reproducibility High: deterministic output under same conditions (parameters) Low: Non deterministic behavior poses challenges to reproducing results Low: As it is difficult to determine when to stop the training for better output Moderate: comparatively reproducible due to better training stability Moderate: comparatively reproducible due to better training stability
Flexibility & Scalability Moderate: tied to assumptions and less adaptable to evolving traffic characteristics High: can be easily extended (e.g. with LSTM) Moderate: conceptually flexible, though practical tuning can be challenging High: versatile in modeling diverse traffic patterns, albeit at a high computational cost High: extremely adaptable across various network scenarios and traffic types
V-E Comparative analysis of different methods
Table VIII provides a summarized comparison of the different approaches used in network traffic generation. Each aspect in the left-most column evaluates key considerations for selecting a generative model:
  • Data Requirements: Measures the amount of real-world data needed for effective training. Models with Low requirements can generate plausible traffic with minimal data, while Very High indicates a reliance on extensive datasets.
  • Model Complexity: Describes the structural and computational complexity of the model, including the number of layers, hyperparameters, and training mechanisms. Low complexity models are simpler and easier to configure, while Very High complexity models will have billions of parameters and require extensive fine-tuning.
  • Training Stability: Evaluates the ease of training a model and the likelihood of instability, such as mode collapse (GANs) or trade-offs in reconstruction (VAEs). More stable models (High) are preferred for reproducibility and reliability.
  • Capture Complex Patterns: Assesses the model’s ability to learn intricate temporal dependencies and non-linear patterns in network traffic. Low models struggle with such details, whereas Very High models excel in both short-term and long-term dependencies.
  • Interpretability: Reflects how transparent the model is in providing insights into its learned representations. Statistical models have high interpretability, whereas deep learning-based models often provide limited intuition.
  • Computational Efficiency: Considers the training and inference costs. High efficiency models are lightweight and computationally inexpensive, whereas Low efficiency models require significant resources.
  • Data Fidelity: Evaluates how closely the generated traffic resembles real-world patterns. High fidelity implies the model generates diverse and realistic traffic flows.
  • Reproducibility: Determines whether the model consistently produces the same results under the same conditions. Low reproducibility models, such as GANs, may produce varying results due to instability in training.
  • Flexibility & Scalability: Examines how well the model adapts to different network conditions and scales with increasing data. Very High scalability models can generalize across multiple network scenarios.
VI Commercially available network traffic simulators
@@
Over the years, traffic generators have predominantly focused on evaluating network protocols and infrastructure rather than the various applications running on top of them. This has been especially true with canonical tools like the venerable Iperf [129], which is often employed to test end-to-end transport protocols, or hardware-based solutions such as IXIA[148], Spirent[37], or XENA[106] which are designed for use cases such as performing RFC 2544 compliant device tests [22]. Traditionally, the division of work between software and hardware-based solutions presented a trade-off, where software tools often fell short in their ability to reach the full core network line rate, a capability more commonly achieved by specialized hardware.
@@
The progression of traffic generators reflects the broader evolution of networking technologies. Modern tools, driven by advancements in software and programmable hardware, offer unprecedented performance and adaptability, enabling more nuanced and application-specific evaluations. As the networking landscape continues to evolve, these tools will undoubtedly play a pivotal role in shaping the methodologies used to ensure the reliability, scalability, and efficiency of future networks. Nonetheless, we can still see nowadays a kind of purpose specialization for each tool where for example Scapy would be used primarily in the context of debugging or proof-of-concept development whereas Spirent would be one of the tools of choice for compliance and system engineering requirement validation.
Table IX: Summary of commercial traffic generators
@@
In order to reflect this de-facto specialization of tools, we present in Table IX an overview of some of the most popular industry-used traffic generators. In particular, we propose to classify those tools under three broad categories: compliance and system engineering, performance evaluation and management, and development and penetration testing. As a result, we can see that due to various regulations, traffic generators for compliance and system engineering tend to be dominated by hardware-based solutions as they made sure to pass numerous certifications for a given piece of equipment. In particular, we note that nowadays, this category is still dominated by tools such as IXIA[148], Spirent[37], or XENA[106].
@@
In the second category of traffic generator applications, the performance and network management, we note the presence of one legacy fully dedicated traffic generator application, Trex [33], and two traffic generator frameworks, SolarWinds and Dynatrace. In the case of Trex, while it started more as a testing tool, its integration with the rest of the CISCO infrastructure, such as the umbrella framework, [^1] https://umbrella.cisco.com/ made this tool use more in a performance and management application with special attention to performance analysis.
@@
Finally, we present in the last identified traffic generator category, namely the development and penetration testing, three tools that are associated with this category. In particular, we mention in Table IX, two tools offering a modular testing capability with an easy-to-use GUI in NetScannTools and Osinato [121] frameworks. Last but not least, we mention a tool that has been gaining more attraction over the years with Scapy, where the user can easily analyze, capture, and replay any type of traffic as long as performance and high throughput are not mandatory due to its python implementation.
VII Open Challenges in Network Traffic Synthesis
While synthetic network traffic data can be a powerful tool for many applications, there remain several challenges that need to be addressed to improve its utility and effectiveness. In this section, we discuss some of the key challenges and technical limitations that complicate the synthesis process.

### v2

@@ IV-C4 Light-weight DMs
Another angle of improvement for DMs were addressed in making the model lightweight such that it can be used for high-quality data generation at edge devices. [97] proposes a DM specifically designed for resource-constrained edge devices. By integrating depthwise separable convolutions into the diffusion process, this model achieves significant reductions in computational overhead while maintaining the diversity and quality of generated malicious traffic. The evaluations demonstrate the ability of the model to produce high-quality synthetic data, improving the training of IDS classifiers even on hardware-limited edge nodes. This lightweight design underscores the potential of diffusion models to address computational challenges, enabling real-time cybersecurity applications at the edge.
IV-C5 Limitations
@@
IV-C6 Summary
The emergence of DMs has revolutionized the field of network traffic analysis, providing a powerful alternative to traditional generative models like GANs. Table VI provides a comprehensive summary of the work done with DMs. Innovative adaptations of DMs, such as DDPM, stable diffusion models, and controlled diffusion models, have demonstrated remarkable capabilities to generate realistic high-quality network traffic data. By effectively modeling the underlying distributions of packet-level features and user interactions with network services, DMs have consistently outperformed previous SOTA GANs in terms of accuracy and realism. The ability of DMs to capture complex temporal and feature correlations while mitigating issues such as mode collapse has solidified their position as the leading generative framework for network traffic synthesis. Further, lightweight diffusion models make the way for high-quality data synthesis in a resource-constrained environment. As research in this area continues to advance, DMs are increasingly pivotal for network security, traffic data analysis, and addressing data scarcity in network environments.
IV-D Transformers
@@
[21] Generic network traffic generation, Attack detection GPT-2 Transport, Network Packet and Flow level No
[115] Generic network traffic generation GPT-2 Transport, Network Packet and Flow level No
[143] Network optimization and traffic flow classification GPT + Linear attention Network Flow level No
[85] Generic network traffic generation, IDS GPT-3 Transport, Network Packet and Flow level No
[91] Cellular network traffic synthesis CPT-GPT Application Event-level sequences No
[180] 3-D Human pose estimation LDT Physical RFID signals No
Transformers were first introduced in 2017 in [169] to address sequence transduction tasks, such as neural machine translation. These tasks involve converting an input sequence into an output sequence. The transformer model is a type of neural network that learns the context of sequential data and generates new data based on its understanding of the relationships between elements within the sequence. Transformers uses attention mechanism to understand context weighing the importance of different elements in a sequence. Table VII summarizes recent work with Transformers
@@
IV-D1 Generative Pretrained Transformers (GPT)
@@
With the introduction of GPT-3 models, authors of [85] leveraged GPT-3 Davinci model and GPT-3 Babbage model to synthesize network traffic for various tasks. The pipeline of this process would be triggered by a user requesting network traffic either using a specific scenario (e.g., normal traffic data, a Ping-of-Death attack, etc.) or by specifying a set of network protocols (e.g., Internet Control Message Protocol [ICMP], Domain Name System [DNS], etc.) Then, a Flow Generator, which is a Python script, produces network flow in text format, i.e. it generates the sequence and textual summary of packets that make up the network traffic, although not creating the actual packets themselves. The flow summary is then passed to the Packet Generator, which is a transformer-based model capable of creating packets using the network flow. The packet generator then creates the specified packets and writes them in .pcap format to a file.GPT-3 Davinci generated training samples for accuracy, while GPT-3 Babbage powered the packet generator for cost-efficient and faster inference.
Similarly, [91] presents CPT-GPT, a high-fidelity traffic generator tailored for cellular network control-plane traffic. CPT-GPT effectively addresses challenges such as stateful semantic correctness, multimodal feature relationships, and long-term data drifts by leveraging a Transformer-based architecture. Through its innovative tokenization scheme and probabilistic output mechanism, CPT-GPT generates realistic traffic while avoiding issues like state violations and mode collapse. Unlike traditional Semi-Markov models (SMMs) that require extensive manual updates to align with evolving the 3rd Generation Partnership Project (3GPP) standards, CPT-GPT achieves better fidelity.
The work in [180] relies on pretrained GPT models and instead introduces a novel framework using latent diffusion transformers (LDT) to generate high-quality, diverse RFID sensing data. Addressing the challenges of collecting paired RFID and 3D human pose data, the synthetic data is used to train a transformer-based kinematics predictor for temporally smooth 3D pose estimation. The framework includes a cross-attention conditioning stage during training and a two-stage velocity alignment design at inference to accurately infer missing joints. This enables the completion of full 25-joint skeletal poses from only 12-joint inputs.
IV-D2 Limitations
@@ IV-D3 Summary
The integration of transformer models, particularly GPT, into network traffic generation marks a major advance in AI-driven data synthesis. Evolving from natural language processing to specialized frameworks like NetGPT and TrafficGPT, these models demonstrate strong versatility in handling complex network traffic data. Advances in tokenization and attention mechanisms address challenges such as flow length limits and class imbalance, improving the quality of synthetic traffic. As they mature, transformer models are expected to transform network analysis, strengthen security, and provide deeper insights through more accurate traffic generation.
V Comparative Analysis of Different Methods
To systematically assess generative approaches for network traffic synthesis, we evaluated each method across nine key aspects using a standardized 1–5 scoring rubric (see Fig. 11). A score of 1 indicates the weakest performance or highest cost/limitation, while 5 reflects ideal performance or efficiency. Where possible, the criteria are grounded in measurable metrics such as number of parameters, required data size, training stability, and computational demands.
Figure 11: Prompt for the AI Tool. The prompt includes the rubric for the evaluation.
We, the authors, selected 86 papers spanning all approaches discussed in this survey and scored them for each aspect using the provided rubrics, drawing on both our expertise and the findings from our literature review. The mean and standard deviation of these scores were then calculated and visualized in Fig. 12 and Fig. 13. To enable readers to quickly obtain scores and analyses for any additional paper they encounter, we provide a prompt (shown below) that can be run on an AI model using the extracted text from the target paper. From our labeled set, we selected random 20 papers as a training set and refined the prompt until it achieved 100% agreement with our manual scores. We then applied the automated script to the remaining 66 papers to verify alignment, achieving 100% accuracy with the final prompt provided in Fig. 11.
For privacy reasons—avoiding the need to upload research papers to cloud-based models—we used GPT-OSS [130] in a local setup. Readers can similarly run the provided prompt on a local model to obtain scores for any synthetic data generation paper. To demonstrate how the rubric applies with the provided prompt, Table VIII presents a model-specific evaluation of NetDiffus [156] using the proposed 9-aspect scoring scheme.
Table VIII: Evaluation of [156] across the 9 scoring aspects
Aspect Score Justification
Data Required 2 Uses approximately 2K samples; strong performance observed with as few as 1K–10K traces.
Model Complexity 3 Implements a 5-layer U-Net with 1K diffusion steps; moderately complex, under 50M parameters.
Training Stability 4 Diffusion training is stable and consistent.
Capture Complex Patterns 4 GASF transformation and iterative denoising allow learning of temporal and nonlinear dependencies.
Interpretability 2 GASF inputs are interpretable, but the diffusion model itself remains largely a black box.
Computational Efficiency 3 Requires GPU acceleration; feasible on a single GPU but not CPU-friendly.
Data Fidelity 4 Synthetic traffic closely matches real traces; achieves significantly lower FID.
Reproducibility 4 DMs yield consistent outputs; high reproducibility across setups is likely.
Flexibility & Scalability 4 Evaluated on diverse traffic types; generalizable with moderate effort.
Figure 12: Comparison of synthetic traffic generation statistical models. Markers: mean scores, shaded region: standard deviation.
Figure 13: Comparison of synthetic traffic generation DL models. Markers: mean scores, shaded region: standard deviation.
V-1 Comparison of Statistical Generative Approaches
To provide a clear understanding of the relative capabilities of various generative approaches for network traffic synthesis, we present a radar plot (Fig. 12) that visualizes key evaluation aspects. As shown in the figure, statistical approaches consistently exhibit strong training stability, interpretability, and reproducibility, which stem from their explicit mathematical formulations and well-understood modeling assumptions. These properties make them well suited for controlled evaluations and analytical studies where transparency and robustness are critical. At the same time, the radar plot reveals clear limitations in capturing complex traffic patterns and achieving high data fidelity, particularly under highly dynamic or non-stationary conditions. Simpler processes such as PP and FBM demonstrate low flexibility & scalability, while more advanced models, including ARIMA and HSM, provide moderate improvements by incorporating additional structure at the cost of increased model complexity. Across the category, computational efficiency remains a distinguishing advantage, enabling practical deployment in resource-constrained environments. Overall, Fig. 12 highlights the strengths of statistical generators in stability and efficiency, alongside their inherent expressiveness limitations.
V-2 Comparison of DL Generative Approaches
Fig. 13 compares representative DL–based traffic generation models using the same evaluation aspects, enabling a direct interpretation of their modeling behavior. Unlike statistical approaches, DL models demonstrate strong performance in capturing complex traffic patterns and achieving high data fidelity, reflecting their ability to learn rich, non-linear representations directly from data. Transformer-based models, in particular, exhibit high flexibility & scalability, driven by attention mechanisms that effectively model long-range dependencies. These gains, however, are accompanied by increased data requirements and model complexity, as evident from the corresponding axes in the radar plot. GANs show comparatively lower training stability, consistent with the challenges associated with mode collapsing. VAEs provide a more stable and computationally efficient alternative but exhibit reduced expressiveness relative to other DL architectures. DMs and Transformers provider higher data fidelity by capturing complex patterns. Across all DL models, interpretability remains limited, underscoring a recurring trade-off between modeling power and transparency. Collectively, Fig. 13 emphasizes the suitability of DL-based generators for data-rich scenarios where realism and scalability are prioritized.
V-3 Statistical vs. DL-Based Synthetic Network Traffic Generation
Building on the family-specific observations in Figs. 12 and 13, Table IX provides a consolidated, high-level comparison between statistical and DL–based traffic generation paradigms. Rather than reporting quantitative scores, the table abstracts the dominant trends observed across both radar plots to highlight paradigm-level trade-offs. Statistical methods consistently emphasize interpretability, data efficiency, training robustness, and computational efficiency due to their explicit formulations and relatively low model complexity. In contrast, DL–based approaches prioritize expressiveness, flexibility & scalability, and data fidelity by learning implicit representations from data, at the expense of transparency and efficiency. This synthesis clarifies that the choice between statistical and DL-based generators is fundamentally application-driven, depending on whether stability and interpretability or expressive power and realism are the primary design objectives.
Table IX: Conceptual comparison of statistical and DL-based synthetic traffic generation models. $\bullet$: strong advantage, $\circ\!\bullet$: mixed or context-dependent, $\circ$: clear limitation.
Aspect Statistical Methods Deep Learning Methods
Data Required $\bullet$ $\circ$
Model Complexity $\bullet$ $\circ$
Training Stability $\bullet$ $\circ\!\bullet$
Capture Complex Patterns $\circ$ $\bullet$
Interpretability $\bullet$ $\circ$
Computational Efficiency $\bullet$ $\circ$
Data Fidelity $\circ\!\bullet$ $\bullet$
Reproducibility $\bullet$ $\circ\!\bullet$
Flexibility & Scalability $\circ$ $\bullet$
V-4 Constraint-Driven Model Selection Matrix
Rather than selecting a generative model based solely on isolated performance metrics, practitioners should align model choice with the dominant constraints of the target application. Drawing on the comparative trends observed in Figs. 12 and 13, we summarize a constraint-driven matrix in Table X that maps common practical requirements to appropriate classes of synthetic network traffic generators.
Table X: Constraint-driven matrix for selecting synthetic network traffic generators.
Primary Constraint Recommended Model Classes and Rationale
Interpretability, data efficiency, analytical transparency Statistical models are preferred due to explicit formulations, low data requirements, stable training, and strong reproducibility. Simple processes (e.g., PP, FBM) suit baseline or stationary traffic, while structured models (e.g., MM, HSM, WD) better capture burstiness and long-range dependence without sacrificing interpretability.
Moderate realism under limited data or compute VAEs provide a balanced trade-off between realism and efficiency, offering stable training, moderate data fidelity, and feasibility on limited computational resources. Suitable when statistical models lack expressiveness but heavier DL models are impractical.
High-fidelity and fine-grained traffic realism DMs are preferred for their strong ability to capture complex temporal and nonlinear dependencies with stable optimization. GANs may be used when faster sampling is required, but their training instability and mode collapse risk must be carefully managed.
Large-scale, heterogeneous, or evolving traffic Transformer-based models are most effective due to attention mechanisms that model long-range dependencies, cross-flow interactions, and domain shifts. Best suited for data-rich scenarios where scalability and flexibility outweigh interpretability concerns.
This matrix highlights that no single generative approach is universally optimal. Increasingly expressive models subsume simpler ones at the cost of higher data requirements, computational overhead, and reduced transparency. Consequently, model selection should be driven by application constraints rather than isolated performance indicators.
VI Commercial network traffic simulators
@@
Traditionally, traffic generators have targeted network protocols and infrastructure, with less emphasis on the applications running above them. This has been especially true with canonical tools like the venerable Iperf [148], which is often employed to test end-to-end transport protocols, or hardware-based solutions such as IXIA[165], Spirent[42], or XENA[122] which are designed for use cases such as performing RFC 2544 compliant device tests [26]. Traditionally, the division of work between software and hardware-based solutions presented a trade-off, where software tools often fell short in their ability to reach the full core network line rate, which is often achieved by specialized hardware.
@@
Table XI: Summary of commercial traffic generators
@@
PROPSIM [83] Hardware Physical layer (Real-world RF channel) Expensive GUI and script RF channel emulation for wireless device testing
CMW series [149] Hardware Physical layer (Wireless protocol signaling) Expensive GUI and script Multi-standard wireless testing and network emulation
@@
Matlab toolbox [166] Software Physical layer Expensive GUI and script System design and waveform generation
GNU Radio [23] Software Physical layer Low/Open source GUI and script Development and testing
QuaDRiGa [73] Software Physical layer Low script 3D radio channel impulse‑response modeling
@@
In order to reflect this de-facto specialization of tools, we present in Table XI an overview of some of the most popular industry-used traffic generators. In particular, we propose to classify those tools under three broad categories: compliance and system engineering, performance evaluation and management, and development and penetration testing. As a result, we can see that due to various regulations, traffic generators for compliance and system engineering tend to be dominated by hardware-based solutions as they made sure to pass numerous certifications for a given piece of equipment. In particular, we note that nowadays, this category is still dominated by tools such as IXIA[165], Spirent[42], or XENA[122].
@@
In the other category of traffic generator applications, the performance and network management, we note the presence of one legacy fully dedicated traffic generator application, Trex [37], and two traffic generator frameworks, SolarWinds and Dynatrace. Considering Trex, while it started more as a testing tool, its integration with the rest of the CISCO infrastructure, such as the umbrella framework(https://umbrella.cisco.com/), made this tool use more in performance and management application with attention to performance analysis.
@@
Then, we present in the last identified traffic generator category, namely the development and penetration testing, three tools that are associated with this category. In particular, we mention in Table XI, two tools offering a modular testing capability with an easy-to-use Graphical User Interface (GUI) in NetScannTools and Osinato [140] frameworks. Also, we mention a tool that has been gaining more attraction over the years with Scapy, where the user can easily analyze, capture, and replay any type of traffic as long as performance and high throughput are not mandatory due to its python implementation.
In addition to traditional packet-level traffic generators, the wireless and physical-layer domain is supported by specialized signal and channel simulators. MATLAB’s wireless toolboxes provide functions for generating standards-compliant RF waveforms and channel models [166]. Open-source Software Defined Radio (SDR) platforms like GNU Radio [23] likewise enable custom wireless signal generation and reception. GNU Radio provides modular signal processing blocks for real-time implementation of Wi-Fi, LTE, or other wireless protocols, and is widely used in both academic and commercial wireless research. Another example is QuaDRiGa [73] (Quasi-Deterministic Radio channel GenerAtor), a MATLAB toolbox for creating realistic 3GPP-compliant MIMO channel impulse responses for system-level simulations. These tools output physical-layer waveforms and measurements (e.g. per-subcarrier channel estimates, RSSI/Reference Signal Received Power[RSRP]) rather than just packet traces, complementing higher-layer traffic generators.
On the hardware side, commercial channel emulators and wireless testers provide realistic over-the-air signal generation. For instance, Keysight’s PROPSIM [83] channel emulation systems can reproduce multipath fading, Doppler shifts, pathloss, and noise in a lab environment to test 5G, Wi‑Fi, and other wireless devices. Similarly, Rohde and Schwarz [149] offers wireless network emulators (e.g. the Rohde & Schwarz [R&S] CMW500/CMX500 series) that generate multi-band LTE/5G/Wi‑Fi signaling with configurable MIMO channel models. These platforms produce actual RF waveforms and report physical layer metrics (RSSI/RSRP, CSI, Error Vector Magnitude [EVM], etc.) under controlled channel conditions. Together, these physical-layer tools extend the scope of traffic generation by embedding real-world wireless channel effects into network testing.
VII Challenges and Future Directions